A car P is 300m behind a car B and both are traveling in the same direction at 25ms–¹. Neglecting the length of the cars.


(i) calculate the acceleration of P.


(ii) it's velocity at the moment of passing B.


(iii) The distance it has travelled in 20s.

1
Expert's answer
2022-03-31T11:48:56-0400

(i)

"v=v_0+at"

"a_B=\frac{\Delta v}{\Delta t}=\frac{27.5-25}{20}=0.125\:\rm m/s^2""d_P=d_B+300"

"v_0t+a_Pt^2\/2=v_0t+a_Bt^2\/2+300"

"a_P*20^2\/2=0.125*20^2\/2+300"

"a_P=1.625\;\rm m/s^2"

(ii)

"v_P=v_0+a_Pt\\
=25+1.625*20=57.5\:\rm m/s"

(iii)

"d_P=v_0t+a_Pt^2/2\\
=25*20+1.625*20^2/2=825\:\rm m"